OpenStudy (anonymous):

The height of a cone-shaped statue is 9 ft, and the diameter is 4 ft. What is the volume of the statue? Use 3.14 to approximate pi, and round answers to two decimal places if necessary. ft3 will give medal for first correct answer and will become fan

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Formula: \(V = \dfrac{1}{3}\pi r^2 h\) Plug in what we know: \(V = \dfrac{1}{3}(3.14) (2)^2 (9)\) Multiply 1/3 * 3.14 * 2 * 2 * 9

OpenStudy (anonymous):

i get 37.68

OpenStudy (igreen):

Correct, that's our answer.

OpenStudy (anonymous):

am I correct @iGreen?

OpenStudy (anonymous):

thank you sooooooo much @iGreen

OpenStudy (igreen):

No problem :)